OpenStudy (anonymous):

Write an equation of a circle with the given center and radius. center (2, 4) and radius 5 (x – 2)2 + (y + 4)2 = 5 (x + 2)2 + (y – 4)2 = 5 (x + 2)2 + (y – 4)2 = 25 (x – 2)2 + (y – 4)2 = 25

OpenStudy (anonymous):

Equation for any circle centre (a,b) radius r is:\[(x-a)^2 + (y-b)^2 = r^2\] So here, the centre is (2,4) and radius 5, so the equation is: \[(x-2)^2 + (y-4)^2 = 25\]